OpenStudy (ahvei):

find the product. 2/3*(-1/2)

OpenStudy (ahvei):

the right answer gets a medal

OpenStudy (misty1212):

\[\frac{2\times (-1)}{3\times 2}\] is the first step then cancel the twos

OpenStudy (ahvei):

i dont think thats the correct answer that not one of my choices @misty1212

OpenStudy (ahvei):



OpenStudy (ahvei):

please help @butterflydreamer and @Smileyside

OpenStudy (anonymous):

Apply this rule: a/b×c/d=ac/bd −2/3×2 Simplify 3×2 to 6 −2/6 Simplify 26 to 13 −1/3

OpenStudy (anonymous):

*2/6 to 1/3

OpenStudy (ahvei):

so whats the answer @Smileyside

OpenStudy (anonymous):

B i think

OpenStudy (ahvei):

thanks @Smileyside

OpenStudy (anonymous):

yup